OpenStudy (anonymous):

Determine whether the graphs of y=-2/3x and 3x-2y=6 are perpendicular lines

OpenStudy (anonymous):

perpendicular lines have opposite reciprocal slopes so both eqn need to be in slope-int form (y=mx+b) to determine slope (m-coefficient of x) in first equation m=-2/3

OpenStudy (anonymous):

3x-2y=6 2y=3x-6 y=3/2 x - 3

OpenStudy (anonymous):

m=-2/3 m=3/2 they are opposite reciprocals (flip numerator and denominator) so yes the lines are perpendicular

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

welcome!